[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Oct 10, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Oct 10 12:19:40 UTC 2017


Here is today's summary of NHC computer operations:
--- PS

----------

TAFB:
1.) None Issues reported since yesterday.
HSU:
1.) Model colors seem fine now. Colors by Intensity option was probably  
checked earlier.

2.) ASCAT data flow issues continue. Problem is upstream but TSB will  
follow it up.
Other items:

NHC RFCs:
1.) RFC for ATCF Memory Upgrade set for Thursday this week.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171010/95aa325c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list